A stock with a beta of 1.3 just paid a dividend of $0.75 and is priced at $42. If the risk-free rate is 3% and the market risk premium is 6%, what is the expected growth rate for the stock? A. 5.02% B. 8.86% C. 12.68% D. 9.01%
Answer #1

Required rate=risk free rate+beta*market risk premium

=3+(1.3*6)=10.8%

Required return=(D1/Current price)+Growth rate

0.108=(0.75*(1+Growth Rate)/42+Growth Rate

0.108=0.017857142+0.017857142*Growth Rate+Growth Rate

Growth Rate=(0.108-0.017857142)/(0.017857142+1)

which is equal to

=8.86%(Approx).